### v1

# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our strong negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further study the impact of different negative sample construction methods and the number of negative samples on the geometric reasoning performance of LMM, yielding fruitful conclusions. The code and dataset are available at https://github.com/THU-KEG/MMGeoLM.
Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
Kai Sun, Yushi Bai, Zhen Yang, Jiajie Zhang, Ji Qi, Lei Hou, Juanzi Li Tsinghua University
11footnotetext: Equal contribution
1 Introduction
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (LMMs) Anthropic (2024); OpenAI (2023); Bai et al. (2023). It requires fine-grained identification of visual elements Lu et al. (2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them He et al. (2024).
[…]
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP Patel et al. (2024); Yang et al. (2023); Goel et al. (2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o OpenAI (2024a), Claude-3 Anthropic (2024), and Qwen2.5-VL Bai et al. (2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
To address these shortcomings, recent research has focused strategies such as fine-tuning on specialized mathematical datasets Gao et al. (2023); Zhang et al. (2024c); Peng et al. (2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs Qi et al. (2024); Wang et al. (2024) by aligning images with corresponding captions. However, the construction of image-caption pairs is inherently susceptible to hallucinations, as many captions are generated by existing LMMs. As a result, training with only these positive samples may lead to spurious alignment Zhang et al. (2024b). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Model Training Training Data MM-MATH
AltCLIP No No 23.8
AltCLIP In-Batch Randomly-sampled Negative (400K) 24.6
AltCLIP MMCLIP ①: Retrieval-based Negative (100K) 26.6
AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
AltCLIP MMCLIP ③: Image Negative (4K) 28.2
AltCLIP MMCLIP ①+② 28.4
AltCLIP MMCLIP ①+②+③ 30.1
Table 1: Comparison of geometric reasoning performance on MM-MATH across LMMs with vision encoders trained with different hard negative data.
In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in the vision encoder. For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives. For image-based contrastive learning, we introduce a novel method that leverages a large language model (LLM) to generate a detailed caption and corresponding diagram generation code for a given geometry problem, forming the positive image sample. The LLM then perturbs the code to create visually similar but geometrically incorrect diagrams, which serve as hard negative samples. Additionally, we modify the original CLIP training framework and propose MMCLIP, a method designed to handle an arbitrary number of hard negative samples centered around a single image or caption. Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
We conduct extensive experiments to evaluate the effectiveness of our CLIP training for LMMs. Specifically, we assess our trained MMGeoLM across four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVISTA, and WE-MATH) and open-ended questions (MM-MATH). Our results show that the proposed model outperforms all existing open-source models on GeoQA and MathVISTA, and achieves state-of-the-art performance on MM-MATH, surpassing GPT-4o by 7.5%. We further analyze the impact of different types of hard negative samples. Notably, we find that negative samples constructed from authentic, exam-based images significantly boost model performance—using just 4K image negatives yields better results than over 100K text negatives. Lastly, we observe that increasing the number of hard negatives improves performance up to a certain threshold, beyond which it leads to diminishing returns and eventual degradation.
Figure 2: Image-based and text-based hard negative construction and the corresponding MMCLIP training method.
2 Method
In this section, we first introduce our proposed geometric diagram strong negative construction method in 2.1. Subsequently, we present the corresponding negative captions approach for geometric elements in 2.2. Finally, we describe a training strategy of handling an arbitrary number of negative examples in 2.3, followed by our newly collected multimodal mathematical dataset in 2.4.
2.1 Negative Images Construction
We construct negative geometric examples by generating geometric diagram codes for positive images and perturbing them to create negatives. Existing generation methods Zhang et al. (2024c); Zou et al. (2024); Wei et al. (2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic exam questions, created by educators using professional geometry software McClintock et al. (2002), produce diagrams that are more diverse and accurate. Specifically, in these authentic geometry problems Sun et al. (2024), geometric elements are often described explicitly (e.g., AB is 8 cm) or implicitly through reasoning (e.g., using the Pythagorean theorem). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs OpenAI (2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional informative annotations (see Appendix A for prompt details).
To handle occasional errors in model-generated code, we adopt a lightweight model GLM et al. (2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro DeepMind (2025), avoiding noise from question-based captions. Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, We design prompts (Appendix A) for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
2.2 Negative Captions Construction
We propose two methods for constructing text-based hard negative samples. The first method retrieves captions similar to each positive example from a large image-caption dataset. The second method generates targeted hard negative samples based on reasoning errors observed during LMM inference. Some examples of these hard negative samples are shown in Figure 1.
Retrieval-based Hard Negative.
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance Karpukhin et al. (2020); Xiong et al. (2020); Huang et al. (2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives Xiong et al. (2020); Zhou et al. (2022); Yang et al. (2024). To address this issue, we build upon the recently proposed MAVIS dataset Zhang et al. (2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS Zhou et al. (2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
Rule-based Hard Negative.
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships Yuksekgonul et al. (2022); Zhang et al. (2024a), while others leverage in-context learning to generate meaningful modifications Patel et al. (2025). However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level without considering downstream reasoning errors. For example, as shown in Figure 1, LMMs incorrectly interpret the parallel line $AB$ and $CD$ as forming triangle $ABC$, leading to errors in solving the geometric problem. Simply altering positive captions without considering LMM reasoning errors provides limited benefits in improving geometric element recognition.
To address this issue, we analyze the evaluation results of LMMs on MM-MATH Sun et al. (2024) and identify four major types of image element recognition errors in their responses. Based on these error types, we design corresponding rule-based strategies to construct hard negative samples:
  1. 1.
    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ are invalid).
  2. 2.
    Shape attributes: Altering properties such as changing squares to rectangles or right triangles to isosceles triangles.
  3. 3.
    Geometric relationships: Modifying relationships such as parallelism between two lines or similarity between two triangles.
  4. 4.
    Numerical values: Adjusting numerical values in captions, such as modifying angles or segment lengths.
We use GLM-4 GLM et al. (2024) to generate 10 rule-based hard negatives for each positive example. An example is shown in Figure 2.
2.3 Hard Negative CLIP Training
Though CLIP-trained vision encoders are widely used in LMMs Radford et al. (2021); Doveh et al. (2023b), they typically adopt in-batch negative sampling during training, which limits their capacity for fine-grained image understanding. While methods like NegCLIP Yuksekgonul et al. (2022) introduce hard negatives, they still rely heavily on in-batch samples. Considering the methods in Section 2.2 have allowed us to construct a large number of hard negatives for each image, we propose MMCLIP training strategy, which focuses on each single image and only uses its corresponding hard negatives for training, entirely abandoning traditional in-batch negatives. In this way, the vision encoder is forced to distinguish subtle differences in all cases, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive example and $T_{j}^{-}$ is the negative samples conducted around the image, using the methods in Section 2.2. Our loss function is formulated as
$\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings of the input instances. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ are expected to be higher than traditional in-batch negatives. From the above equation, we calculate the gradient for positive and negative samples:
$\frac{\partial\mathcal{L}}{\partial s_{i}^{+}}=-\frac{\sum_{j=1}^{N}\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (2)
$\frac{\partial\mathcal{L}}{\partial s^{-}_{j}}=\frac{\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (3)
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. We further validate the effectiveness of hard negatives and analyze the impact of their quantity in our experiments in Section 4.2.
2.4 Supervised Fine-tuning Data
[…]
3 Multimodal Mathematical Training
3.1 Architecture
We adopt the LLaVA architecture Liu et al. (2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use Mammoth2-7B Yue et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025). The vision encoder is based on AltCLIP Chen et al. (2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
3.2 Training Details
In the vision-text alignment stage, we first pre-train AltCLIP on the 400K MAVIS image-caption alignment dataset Zhang et al. (2024c) to improve the model’s generalization for geometric problems. Next, we fine-tune the model following the MMCLIP training strategy, using 100K hard caption negative samples, each consisting of 10 rule-based and 30 retrieval-based negatives to enhance the model’s understanding of geometric elements. In addition, we fine-tune the model on 4K hard image negative samples, where each example includes 10 negative geometric diagrams. In the second stage, we tune the MLP adapter using 67K G-LLAVA Gao et al. (2023) image-text alignment data. In the third stage, we perform supervised fine-tuning on a combined dataset comprising 300K MAVIS instruction data Zhang et al. (2024c), 117K G-LLaVA instruction data Gao et al. (2023), and 17K open-ended geometric problems. The open-ended dataset includes 12K problems from our collected dataset and 5K randomly sampled MM-MATH Sun et al. (2024) geometric problems, with the rest 700 problems reserved for evaluation. We name the trained model as MMGeoLM.
4 Experiments
To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs. Additionally, we analyze the impact of different vision encoder types and training strategies, as well as the number of hard negatives in MMGeoLM.
4.1 LMM Geometric Reasoning Evaluation
4.1.1 Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA Chen et al. (2021), a geometric question-answering task based on plane geometry; We-Math Qiao et al. (2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista Lu et al. (2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-MATH Sun et al. (2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-MATH, we evaluate models across easy, medium, and hard categories using a test set of 700 problems excluded from training. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
4.1.2 Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus Anthropic (2024), Claude-3.5-Sonnet, GPT-4o OpenAI (2024a), and GPT-4V. Open-source LMMs: MAVIS-7B Zhang et al. (2024c), G-LLaVA-7B Gao et al. (2023), InternVL2-8B Chen et al. (2024), LLaVA-OneVision-7B Li et al. (2024), Qwen2.5-VL-7B-Instruct Bai et al. (2025), Chimera-Reasoner-8B Peng et al. (2024b), InternLM-XComposer2-7B Dong et al. (2024), Phi-3-Vision-128K-Instruct Abdin et al. (2024), and Math-LLaVA-13B Shi et al. (2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks Chen et al. (2021); Lu et al. (2023) since 2023. Additionally, we include human evaluation baselines, using scores from prior studies Zhang et al. (2024c); Sun et al. (2024).
Model GeoQA MathVista We-Math MM-MATH
GPS ALG S2 S3 Easy Med. Hard Avg
Human* 92.3 48.4 50.9 - - 90.7 81.9 47.6 80.4
Claude-3-Opus* 44.5 46.3 46.6 32.9 23.0 29.5 19.3 3.6 20.3
Claude-3.5-Sonnet 65.1 66.3 68.4 64.7 62.1 34.4 31.9 13.6 31.7
GPT-4V* - 50.5 53.0 49.2 38.2 37.8 21.2 1.8 23.1
GPT-4o 58.9 62.7 65.4 58.0 43.6 45.8 30.0 10.9 31.8
MAVIS-7B* 68.3 64.1 59.2 37.9 34.6 - - - -
G-LLaVA-7B* 67.0 56.7 - 30.1 32.7 - - - -
Math-LLaVA-13B 48.4 55.6 55.0 31.7 23.0 - - - -
InternVL2-8B 56.4 60.5 60.8 41.1 37.1 33.6 21.7 9.0 23.3
LLaVA-OneVision-7B 64.6 54.6 53.1 28.7 22.3 40.5 24.8 10.5 27.0
Chimera-Reasoner-8B 69.6 48.5 42.6 29.9 31.6 36.2 22.2 9.0 24.1
Phi-3-Vision-128K-Instruct 29.5 38.8 40.0 33.3 30.1 12.9 7.1 0.0 7.8
Qwen2.5-VL-7B-Instruct 59.0 67.7 66.9 58.1 50.6 50.7 32.2 14.2 34.8
InternLM-XComposer2-7B 38.8 63.0 56.6 33.1 33.0 18.9 12.2 4.5 13.1
MMGeoLM-Mammoth2-7B 68.5 68.7 69.5 41.5 37.5 52.5 32.4 4.5 36.9
   w/o MMCLIP 55.4 52.2 51.3 34.1 35.3 42.5 21.8 4.5 26.8
   Original AltCLIP 46.7 45.8 46.8 30.1 29.2 38.2 20.4 1.6 25.9
MMGeoLM-Qwen2.5-7B 69.2 69.8 68.1 39.8 36.4 55.3 36.9 9.0 39.2
   w/o MMCLIP 56.3 54.1 50.3 35.2 35.1 50.5 28.8 4.5 34.3
   Original AltCLIP 50.7 47.6 48.9 33.1 32.7 52.5 30.8 4.5 33.7
Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and MM-MATH. ‘S2/S3’ denote the two-step settings in We-Math. Results marked * are taken from their original papers.
4.1.3 Overall Results
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-MATH benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA Peng et al. (2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-Mammoth2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-Mammoth2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-MATH benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the original AltCLIP and without MMCLIP (AltCLIP trained only with in-batch learning on MAVIS), the vision encoder trained with our constructed hard negative samples and the MMCLIP training method significantly enhances MMGeoLM’s geometric understanding capabilities.
Figure 3: Performance evaluation of MMGeoLMs built with six AltCLIP trained on different contrastive datasets.
Figure 4: Comparison of two image negative strategies: Mavis uses 100K samples from the Mavis dataset, while MM-MATH-Align uses our 4K constructed samples.
Figure 5: Retrieval performance of AltCLIP variants trained with different strategies. ‘Random + X’ means sequential training with random negatives followed by negatives created with the X strategy.
4.2 Experiments on Vision Encoder
We further quantitatively evaluate the performance of various vision encoders trained with different datasets and strategies. The experiments are divided into three parts: (1) performance evaluation of six LMMs trained with different AltCLIP variants across two benchmarks, (2) top-1 positive example retrieval performance evaluation of six AltCLIP models, and (3) varying number of hard caption negatives.
4.2.1 Setup
We compare the following six types of AltCLIP models are as follows:
  • Original AltCLIP: No additional data training.
  • Random AltCLIP: Trained with randomly selected negative samples within a batch (batch size = 10) using 400K MAVIS data.
  • Retrieved AltCLIP: Trained with 100K samples, each with 10 retrieval-based caption negatives.
  • Rule-based AltCLIP: Trained with 100K samples, each with 10 rule-based caption negatives.
  • Negative-images AltCLIP: Trained with 4K samples, each with 10 hard image negatives.
  • Retrieved + Rule-based + Negative-images: Trained on all three datasets combined.
The Random AltCLIP follows OpenAI’s training approach Radford et al. (2021), while the Retrieved, Rule-based, Negatives-images and Retrieved + Rule-based + Negatives-images AltCLIP models employ our proposed MMCLIP training approach. In the second and third stages, LMMs are trained on 67K G-LLaVA alignment data and 17K open-ended geometric reasoning problems.
For the retrieval experiments, we assess the retrieval performance of trained AltCLIP models by selecting positive samples from a large pool of negative samples. Retrieval performance is evaluated using the Hit@1 metric. We construct four types of negative datasets: Random Negative: Caption negatives randomly selected in a training batch. Retrieval Negative: Caption negatives retrieved using the SimANS Zhou et al. (2022) model, selecting the top 100 most similar captions to the positive sample. Rule-based Negative: Caption negatives generated according to the rules defined in section 2.2. Image Negative: Image negatives generated from real-world geometric problems by MM-MATH.
For the number of negative experiments, we investigate the impact of hard caption negative samples on MMGeoLM varying from 5 to 50.
4.2.2 Results
As shown in Figure 4, LMMs trained with hard negative samples consistently improve performance on the GeoQA and MM-MATH benchmarks. Notably, on MM-MATH, Negative-images AltCLIP, trained with only 4K image negative samples, surpasses Retrieved AltCLIP and Rule-based AltCLIP by 0.2% and 1.7%, respectively. On MM-MATH, Rule-based, Retrieved, and Image-negative AltCLIP consistently outperform Random AltCLIP on easy and medium-level questions. These results confirm the effectiveness of both hard negative construction methods. Furthermore, AltCLIP trained with a combination of caption and image negatives achieves the best performance on both GeoQA and MM-MATH, demonstrating their complementary effect in improving model performance. Further experiments on image negatives (Figure 4) reveal that Negative-image-MM-MATH-Align, constructed from real-world images, outperforms Negative-image-Mavis. This suggests that high-quality, realistic image negatives have a more pronounced impact on model performance than large-scale synthetic counterparts.
The results of the retrieval performance experiments are shown in Figure 5. Under the negative-images setting, Negative-images AltCLIP achieves a retrieval score of 23%, while other AltCLIP variants remain below 15%. These results highlight the difficulty of retrieving realistic images. Notably, the dataset includes only 4,021 image negative samples, which is likely insufficient to fully exploit the model’s performance potential. In the retrieval-based and rule-based negative settings, Random AltCLIP outperforms Rule-based AltCLIP in retrieval-based settings and exceeds Retrieved AltCLIP in rule-based settings. These results indicate that while hard negative training enhances performance within specific domains, it may limit generalization. Furthermore, pre-training on random negatives followed by hard negatives improves both generalization and domain-specific performance, ultimately enhancing overall model effectiveness.
Figure 6: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative samples.
As shown in Figure 6, MMGeoLM’s performance on MM-MATH improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point. These results indicate that hard caption negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance. Training with only five hard negative samples outperforms training with ten randomly selected negatives (as shown in Table 1), demonstrating that hard negatives enhance multimodal mathematical reasoning more effectively.
5 Related Work
Mathematics-related research has recently received significant attention in LLMs and LMMs. For text-only mathematical reasoning, several works employ external tools like Tora Gou et al. (2023) or intermediate step decomposition methods such as MAmmoTH Yue et al. (2023), Metamath Yu et al. (2023) and Math-Shepherd Yue et al. (2023). For multimodal mathematics tasks involving images, benchmarks have been proposed, including mathematical competition-oriented OlympiadBench He et al. (2024), geometry-focused datasets like Geometry3K Lu et al. (2021), VisScience Jiang et al. (2024), UniGeo Chen et al. (2022a) and GeoQA Chen et al. (2022a). Many works take effort on training multimodel math, including G-llava Gao et al. (2023), Meta-LLaVA Shi et al. (2024) and MAVIS Zhang et al. (2024c). However, these multimodal approaches typically suffer from inaccuracies in recognizing geometric elements, where our paper specifically addresses the enhancement of visual recognition accuracy.
Vision encoders play a crucial role in understanding image content. While previous works Zhang et al. (2024a); Doveh et al. (2023a, b); Singh et al. (2023) have leveraged negative captions to enhance image comprehension, but not address the recognition of geometric elements within images. Methods such as NegCLIP Yuksekgonul et al. (2022) and TriCLIP Yang et al. (2024) focus on learning from negative samples, but their applicability is limited by the constrained number of negative samples. In this work, we propose a robust negative learning strategy that explicitly constructs negative samples for geometric elements while enabling arbitrary expansion, thereby improving geometric understanding in multimodal tasks.
6 Conclusion
This paper introduced MMCLIP, a novel hard negative contrastive learning framework to enhance LMMs’ geometric reasoning by refining their vision encoders with systematically generated image and text-based negatives. Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks. This work demonstrates the power of tailored hard negative sampling for developing LMMs with fine-grained visual understanding in specialized domains such as geometric understanding and reasoning.
Limitations
Despite the effectiveness of the proposed image hard negatives constructed by perturbing Python scripts, the method heavily relies on the accuracy of LLM-generated code and captions. Any systematic biases in these synthetic constructions may introduce artifacts that differ from human-designed geometric problems. As a result, the model’s performance on real-world visual inputs remains uncertain and requires further validation on diverse real-world datasets.
Appendix A Prompts
After multiple experimental iterations, we constructed two different types of prompts for our experiments, as illustrated in Figures 7 and 8. Specifically, Figure 7 demonstrates the prompt constructed by inputting the textual description and answer of realistic geometry problems to obtain the corresponding geometric diagram Python script. Meanwhile, Figure 8 illustrates the prompt obtained by modifying captions from positive examples to create slightly perturbed negative samples.
Figure 7: Python script prompt from question and true answer in geometric problems
Figure 8: Negative caption generation prompt from positive captions
Appendix B Figure Comparison
Figure 9 presents illustrative examples of image negative sample, covering analytical geometry and planar geometry. Geometric elements include triangles, quadrilaterals, as well as various relational properties such as perpendicularity and intersection. Additionally, certain metric properties absent from the original diagrams are explicitly indicated. Specifically, the first column shows the original diagrams, and the second column depicts diagrams generated from the given questions and the true answer, which, despite slight discrepancies, accurately capture the overall outlines of the original diagrams. The third and fourth columns display the constructed negative diagram instances—generated based upon their corresponding positive examples—demonstrating notable similarities but also clear and meaningful distinctions.
Figure 9: Comparison of generated geometric figures. The first column shows the original images, the second column presents the positive samples, and the last two columns illustrate the constructed negative samples.
Appendix C Case Study
Figure 10: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. MMGeoLM correctly answers the problem using a different solution compared to True Answer, while GPT-4o generates a non-existent element $\triangle ABD$.
Figure 11: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
C.1 Case Study
In Figures 10 and 11, we compare the solutions generated by GPT-4o and our MMGeoLM. In Figure 10, the red text highlights GPT-4o’s misrecognition of two triangles in the image, leading to an incorrect final result. In contrast, MMGeoLM produces a correct solution, albeit through a reasoning process different from the ground-truth answer. This highlights MMGeoLM’s capability to generate diverse problem-solving approaches. Figure 11 illustrates errors in image-based reasoning for both MMGeoLM and GPT-4o. While both models make mistakes, MMGeoLM’s reasoning aligns more closely with the correct solution, demonstrating improved mathematical reasoning. Overall, our trained MMGeoLM model demonstrates greater reasoning diversity and enhanced geometric problem-solving performance compared to current mainstream models.

### v2

@@
 To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity.
-We train CLIP using our strong negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving.
+We train CLIP using our hard negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving.
 Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks.
 Even with a size of 7B, it can rival powerful closed-source models like GPT-4o.
-We further study the impact of different negative sample construction methods and the number of negative samples on the geometric reasoning performance of LMM, yielding fruitful conclusions.
-The code and dataset are available at https://github.com/THU-KEG/MMGeoLM.
-Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
-Kai Sun, Yushi Bai, Zhen Yang, Jiajie Zhang, Ji Qi, Lei Hou, Juanzi Li Tsinghua University
-11footnotetext: Equal contribution
-1 Introduction
-Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (LMMs) Anthropic (2024); OpenAI (2023); Bai et al.
-(2023).
-It requires fine-grained identification of visual elements Lu et al.
-(2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them He et al.
-(2024).
+We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations.
+These analyses yield important insights into optimizing hard negative strategies for geometric reasoning tasks.
+https://github.com/THU-KEG/MMGeoLM.
+Introduction
+Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic 2024; OpenAI 2023; Bai et al.
+2023).
+It requires fine-grained identification of visual elements (Lu et al.
+2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al.
+2024).
+However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al.
+2024; Yang et al.
+2023; Goel et al.
+2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning.
+Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers.
+As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI 2024a), Claude-3 (Anthropic 2024), and Qwen2.5-VL (Bai et al.
+2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
 […]
-However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP Patel et al.
-(2024); Yang et al.
-(2023); Goel et al.
-(2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning.
-Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers.
-As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o OpenAI (2024a), Claude-3 Anthropic (2024), and Qwen2.5-VL Bai et al.
-(2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
-To address these shortcomings, recent research has focused strategies such as fine-tuning on specialized mathematical datasets Gao et al.
-(2023); Zhang et al.
-(2024c); Peng et al.
-(2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs Qi et al.
-(2024); Wang et al.
-(2024) by aligning images with corresponding captions.
-However, the construction of image-caption pairs is inherently susceptible to hallucinations, as many captions are generated by existing LMMs.
-As a result, training with only these positive samples may lead to spurious alignment Zhang et al.
-(2024b).
+To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al.
+2023; Zhang et al.
+2024b; Peng et al.
+2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al.
+2024; Wang et al.
+2024) by aligning images with corresponding captions.
+However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al.
+2024a; Doveh et al.
+2023a).
 To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations.
 Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
-Model Training Training Data MM-MATH
+Model Training Training Data $M$-Math
 AltCLIP No No 23.8
@@
 AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
-AltCLIP MMCLIP ③: Image Negative (4K) 28.2
+AltCLIP MMCLIP ③: Image-based Negative (4K) 28.2
 AltCLIP MMCLIP ①+② 28.4
@@
 Table 1: Comparison of geometric reasoning performance on MM-Math across LMMs with vision encoders trained with different hard negative data.
-In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in the vision encoder.
+In this work, we propose two types of hard negative sample construction methods, i.e., image-based and text-based, to enhance fine-grained geometric element recognition in vision encoder.
 For text-based contrastive learning, we design two strategies to create negative captions for a geometry image: (1) a retrieval-based method that employs dense retrieval in a geometric-domain text corpus to select captions with high lexical similarity but differing content as negative samples; and (2) a rule-based method that modifies key geometric attributes in the captions, such as shapes, angles, and lengths, while keeping other elements unchanged, thereby producing negative samples that bear similar appearance but with key distinct information from the positives.
@@
 Table 1 shows the effectiveness of MMCLIP training on different hard negative sets.
-We conduct extensive experiments to evaluate the effectiveness of our CLIP training for LMMs.
-Specifically, we assess our trained MMGeoLM across four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVISTA, and WE-MATH) and open-ended questions (MM-MATH).
-Our results show that the proposed model outperforms all existing open-source models on GeoQA and MathVISTA, and achieves state-of-the-art performance on MM-MATH, surpassing GPT-4o by 7.5%.
-We further analyze the impact of different types of hard negative samples.
-Notably, we find that negative samples constructed from authentic, exam-based images significantly boost model performance—using just 4K image negatives yields better results than over 100K text negatives.
-Lastly, we observe that increasing the number of hard negatives improves performance up to a certain threshold, beyond which it leads to diminishing returns and eventual degradation.
+We evaluate our trained MMGeoLM on four geometric benchmarks spanning two categories: choice-based questions (GeoQA, MathVista, and We-Math) and open-ended questions (MM-Math).
+Results show that the proposed model outperforms all existing open-source models on GeoQA and MathVista, and achieves state-of-the-art performance on MM-Math, surpassing GPT-4o by 7.4%.
+Ablation studies confirm both text-based and image-based hard negatives benefit model performance.
+Notably, using only 4K image-based negatives yields better results than 100K retrieval-based ones.
+These findings demonstrate that our hard negative construction and training strategy significantly enhances reasoning accuracy in geometric reasoning.
 Figure 2: Image-based and text-based hard negative construction and the corresponding MMCLIP training method.
-2 Method
-In this section, we first introduce our proposed geometric diagram strong negative construction method in 2.1.
-Subsequently, we present the corresponding negative captions approach for geometric elements in 2.2.
-Finally, we describe a training strategy of handling an arbitrary number of negative examples in 2.3, followed by our newly collected multimodal mathematical dataset in 2.4.
-2.1 Negative Images Construction
-We construct negative geometric examples by generating geometric diagram codes for positive images and perturbing them to create negatives.
-Existing generation methods Zhang et al.
-(2024c); Zou et al.
-(2024); Wei et al.
-(2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems.
-In contrast, authentic exam questions, created by educators using professional geometry software McClintock et al.
-(2002), produce diagrams that are more diverse and accurate.
-Specifically, in these authentic geometry problems Sun et al.
-(2024), geometric elements are often described explicitly (e.g., AB is 8 cm) or implicitly through reasoning (e.g., using the Pythagorean theorem).
-Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs OpenAI (2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images.
+Method
+Negative Images Construction
+We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives.
+Existing generation methods (Zhang et al.
+2024b; Zou et al.
+2024; Wei et al.
+2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems.
+In contrast, authentic geometry problems (Sun et al.
+2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock, Jiang, and July 2002).
+Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images.
 As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption.
-Executing the code produces diagrams that not only replicate the original figures but may also include additional informative annotations (see Appendix A for prompt details).
-To handle occasional errors in model-generated code, we adopt a lightweight model GLM et al.
-(2024) to correct syntax issues when the code fails.
-Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro DeepMind (2025), avoiding noise from question-based captions.
+Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
+To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al.
+2024) to correct syntax issues when the code fails.
+Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind 2025).
 Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
-To create the negative images, We design prompts (Appendix A) for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically.
+To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically.
 We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions.
 These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
-2.2 Negative Captions Construction
+Negative Captions Construction
 We propose two methods for constructing text-based hard negative samples.
 The first method retrieves captions similar to each positive example from a large image-caption dataset.
-The second method generates targeted hard negative samples based on reasoning errors observed during LMM inference.
-Some examples of these hard negative samples are shown in Figure 1.
+The second method generates targeted hard negative samples based on reasoning errors observed during LMMs’ inference.
 Retrieval-based Hard Negative.
-Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance Karpukhin et al.
-(2020); Xiong et al.
-(2020); Huang et al.
-(2020).
-However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives Xiong et al.
-(2020); Zhou et al.
-(2022); Yang et al.
-(2024).
-To address this issue, we build upon the recently proposed MAVIS dataset Zhang et al.
-(2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption.
-We use SimANS Zhou et al.
-(2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
+Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al.
+2020; Xiong et al.
+2020; Huang et al.
+2020).
+However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al.
+2020; Zhou et al.
+2022; Yang et al.
+2024).
+To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al.
+2024b), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption.
+We use SimANS (Zhou et al.
+2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
 Rule-based Hard Negative.
-Existing studies modify positive captions by randomly altering nouns, attributes, and relationships Yuksekgonul et al.
-(2022); Zhang et al.
-(2024a), while others leverage in-context learning to generate meaningful modifications Patel et al.
-(2025).
-However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level without considering downstream reasoning errors.
-For example, as shown in Figure 1, LMMs incorrectly interpret the parallel line $AB$ and $CD$ as forming triangle $ABC$, leading to errors in solving the geometric problem.
-Simply altering positive captions without considering LMM reasoning errors provides limited benefits in improving geometric element recognition.
-To address this issue, we analyze the evaluation results of LMMs on MM-MATH Sun et al.
-(2024) and identify four major types of image element recognition errors in their responses.
-Based on these error types, we design corresponding rule-based strategies to construct hard negative samples:
+Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al.
+2022; Zhang, Awal, and Agrawal 2024), while others leverage in-context learning to generate meaningful modifications (Patel et al.
+2025) .
+However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
+To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al. 2024) and identify four major types of image element recognition errors:
   1. 1.
-    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ are invalid).
+    Geometric element ordering: Modifying the sequence of alphabetical order in geometric diagrams while ensuring the new order does not match the original cyclically (e.g., $ABCD$ changing to $CDAB$ is invalid).
   2. 2.
@@
     Numerical values: Adjusting numerical values in captions, such as modifying angles or segment lengths.
-We use GLM-4 GLM et al. (2024) to generate 10 rule-based hard negatives for each positive example. An example is shown in Figure 2.
-2.3 Hard Negative CLIP Training
-Though CLIP-trained vision encoders are widely used in LMMs Radford et al.
-(2021); Doveh et al.
-(2023b), they typically adopt in-batch negative sampling during training, which limits their capacity for fine-grained image understanding.
-While methods like NegCLIP Yuksekgonul et al.
-(2022) introduce hard negatives, they still rely heavily on in-batch samples.
-Considering the methods in Section 2.2 have allowed us to construct a large number of hard negatives for each image, we propose MMCLIP training strategy, which focuses on each single image and only uses its corresponding hard negatives for training, entirely abandoning traditional in-batch negatives.
-In this way, the vision encoder is forced to distinguish subtle differences in all cases, thereby enhancing its fine-grained geometry understanding capacities.
-Formally, given a training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive example and $T_{j}^{-}$ is the negative samples conducted around the image, using the methods in Section 2.2.
-Our loss function is formulated as
+To address these error types, we design rule-based strategies that use GLM-4 (GLM et al. 2024) to modify each positive caption and generate 10 hard negative examples.
+These text-based negatives are used in model training, and two representative types are shown in Figure 2.
+Hard Negative CLIP Training
+Figure 3: Overview of the MMGeoLM training pipeline, including main and each ablation experiment configurations, training strategies, and datasets.
+Though CLIP-trained vision encoders are widely used in LMMs (Radford et al.
+2021; Doveh et al.
+2023b), they typically combine random examples for a batch, which limits their capacity for fine-grained image understanding during training.
+While methods like NegCLIP (Yuksekgonul et al.
+2022) introduce hard negatives, they still rely heavily on randomly selected negatives within a batch.
+Considering the methods in the above section have allowed us to construct a large number of hard negatives, we propose a MMCLIP training strategy, which focuses on each single sample and only uses its corresponding hard negatives for training.
+In this way, the vision encoder is forced to distinguish subtle differences, thereby enhancing its fine-grained geometry understanding capacities.
+Formally, given a text-based negative training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive samples and $T_{j}^{-}$ is the negative samples conducted around the image.
+Our loss function is formulated as follows:
 $\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
-where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings of the input instances.
-Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ are expected to be higher than traditional in-batch negatives.
+where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings.
+Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ is expected to be higher than a randomly chosen one.
 From the above equation, we calculate the gradient for positive and negative samples:
@@
 The training strategy is illustrated in Figure 2.
-We further validate the effectiveness of hard negatives and analyze the impact of their quantity in our experiments in Section 4.2.
-2.4 Supervised Fine-tuning Data
+Unlike in-batch training, which relies on batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for rule-based or retrieval-based unimodal negatives.
+We further validate different training strategies only for hard negatives in the ablation experiments.
+Supervised Fine-tuning Data
 […]
-3 Multimodal Mathematical Training
-3.1 Architecture
-We adopt the LLaVA architecture Liu et al.
-(2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone.
-For the LLM backbone, we use Mammoth2-7B Yue et al.
-(2024) and Qwen2.5-7B-Instruct Qwen et al.
-(2025).
-The vision encoder is based on AltCLIP Chen et al.
-(2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
-3.2 Training Details
-In the vision-text alignment stage, we first pre-train AltCLIP on the 400K MAVIS image-caption alignment dataset Zhang et al.
-(2024c) to improve the model’s generalization for geometric problems.
-Next, we fine-tune the model following the MMCLIP training strategy, using 100K hard caption negative samples, each consisting of 10 rule-based and 30 retrieval-based negatives to enhance the model’s understanding of geometric elements.
-In addition, we fine-tune the model on 4K hard image negative samples, where each example includes 10 negative geometric diagrams.
-In the second stage, we tune the MLP adapter using 67K G-LLAVA Gao et al.
-(2023) image-text alignment data.
-In the third stage, we perform supervised fine-tuning on a combined dataset comprising 300K MAVIS instruction data Zhang et al.
-(2024c), 117K G-LLaVA instruction data Gao et al.
-(2023), and 17K open-ended geometric problems.
-The open-ended dataset includes 12K problems from our collected dataset and 5K randomly sampled MM-MATH Sun et al.
-(2024) geometric problems, with the rest 700 problems reserved for evaluation.
-We name the trained model as MMGeoLM.
-4 Experiments
-To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs.
-Additionally, we analyze the impact of different vision encoder types and training strategies, as well as the number of hard negatives in MMGeoLM.
-4.1 LMM Geometric Reasoning Evaluation
-4.1.1 Datasets and Metrics
-We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended.
-The multiple-choice benchmark includes GeoQA Chen et al.
-(2021), a geometric question-answering task based on plane geometry; We-Math Qiao et al.
-(2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista Lu et al.
-(2023), widely used for evaluating LMMs’ performance.
-The open-ended benchmark is MM-MATH Sun et al.
-(2024), which features high discriminative difficulty sourced from secondary school-level problems.
-We use accuracy (ACC) as the metric.
-For We-Math, we assess multi-step problems (S2 and S3), while for MM-MATH, we evaluate models across easy, medium, and hard categories using a test set of 700 problems excluded from training.
-For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
-4.1.2 Baselines
-The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus Anthropic (2024), Claude-3.5-Sonnet, GPT-4o OpenAI (2024a), and GPT-4V.
-Open-source LMMs: MAVIS-7B Zhang et al.
-(2024c), G-LLaVA-7B Gao et al.
-(2023), InternVL2-8B Chen et al.
-(2024), LLaVA-OneVision-7B Li et al.
-(2024), Qwen2.5-VL-7B-Instruct Bai et al.
-(2025), Chimera-Reasoner-8B Peng et al.
-(2024b), InternLM-XComposer2-7B Dong et al.
-(2024), Phi-3-Vision-128K-Instruct Abdin et al.
-(2024), and Math-LLaVA-13B Shi et al.
-(2024).
-For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks Chen et al.
-(2021); Lu et al.
-(2023) since 2023.
-Additionally, we include human evaluation baselines, using scores from prior studies Zhang et al.
-(2024c); Sun et al.
-(2024).
-Model GeoQA MathVista We-Math MM-MATH
-GPS ALG S2 S3 Easy Med. Hard Avg
+Multimodal Mathematical Training
+Architecture
+We adopt the LLaVA architecture (Liu et al.
+2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone.
+For the LLM backbone, we use MAmmoTH2-7B (Yue et al.
+2024) and Qwen2.5-7B-Instruct (Qwen et al.
+2025).
+The vision encoder is based on AltCLIP (Chen et al.
+2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
+Training Details
+Model GeoQA MathVista We-Math $M$-Math
+GPS ALG S2 S3 Easy Med Hard Avg
 Human* 92.3 48.4 50.9 - - 90.7 81.9 47.6 80.4
@@
   Original AltCLIP 50.7 47.6 48.9 33.1 32.7 52.5 30.8 4.5 33.7
-Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and MM-MATH. ‘S2/S3’ denote the two-step settings in We-Math. Results marked * are taken from their original papers.
-4.1.3 Overall Results
+Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and $M$-Math. ‘S2/S3’ denotes the two-step settings in We-Math. Results marked * are taken from their original papers.
+Our training pipeline, illustrated in Figure 3, consists of three stages covering both main and ablation experiments.
+In the first stage, we perform vision-text alignment training.
+For text-based hard negatives, we use the Mavis image-caption alignment dataset (Zhang et al.
+2024b).
+For image-based negatives, we collect 4K geometry questions from real middle-school exams.
+Additional details, including the types of negatives and the negative-to-positive ratio, are provided in Figure 3.
+In the second stage, we tune the MLP adapter using 67K image-text pairs from the G-LLaVA dataset (Gao et al.
+2023).
+In the third stage, we conduct supervised fine-tuning.
+For main experiments, we use a combined dataset comprising 300K Mavis instruction data (Zhang et al.
+2024b), 117K G-LLaVA instruction data (Gao et al.
+2023), and 17K open-ended geometry questions.
+For ablation studies, we only use the 17K newly collected geometry problems for efficient validation.
+We name the final trained model as MMGeoLM.
+Experiments
+To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs.
+Additionally, we analyze the impact of different vision encoder types, image-based negatives and training strategies in the subsequent ablation studies.
+LMM Geometric Reasoning Evaluation
+Datasets and Metrics
+We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended.
+The multiple-choice benchmark includes GeoQA (Chen et al.
+2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al.
+2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al.
+2023), widely used for evaluating LMMs’ performance.
+The open-ended benchmark is MM-Math (Sun et al.
+2024), which features high discriminative difficulty sourced from secondary school-level problems.
+We use accuracy (ACC) as the metric.
+For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems.
+For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
+Baselines
+The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic 2024), Claude-3.5-Sonnet, GPT-4o (OpenAI 2024a), and GPT-4V.
+Open-source LMMs: MAVIS-7B (Zhang et al.
+2024b), G-LLaVA-7B (Gao et al.
+2023), InternVL2-8B (Chen et al.
+2024), LLaVA-OneVision-7B (Li et al.
+2024), Qwen2.5-VL-7B-Instruct (Bai et al.
+2025), Chimera-Reasoner-8B (Peng et al.
+2024b), InternLM-XComposer2-7B (Dong et al.
+2024), Phi-3-Vision-128K-Instruct (Abdin et al.
+2024), and Math-LLaVA-13B (Shi et al.
+2024).
+For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al.
+2021; Lu et al.
+2023).
+Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al.
+2024b; Sun et al.
+2024).
+Overall Results
 As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks.
 On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B.
-As Chimera-Reasoner-8B was trained on GeoQA Peng et al.
-(2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset.
+As Chimera-Reasoner-8B was trained on GeoQA (Peng et al.
+2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset.
 The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach.
@@
 In contrast, hard problems involve multi-step reasoning, where its impact is more limited.
-Compared to using the original AltCLIP and without MMCLIP (AltCLIP trained only with in-batch learning on MAVIS), the vision encoder trained with our constructed hard negative samples and the MMCLIP training method significantly enhances MMGeoLM’s geometric understanding capabilities.
-Figure 3: Performance evaluation of MMGeoLMs built with six AltCLIP trained on different contrastive datasets.
-Figure 4: Comparison of two image negative strategies: Mavis uses 100K samples from the Mavis dataset, while MM-MATH-Align uses our 4K constructed samples.
-Figure 5: Retrieval performance of AltCLIP variants trained with different strategies. ‘Random + X’ means sequential training with random negatives followed by negatives created with the X strategy.
-4.2 Experiments on Vision Encoder
-We further quantitatively evaluate the performance of various vision encoders trained with different datasets and strategies.
-The experiments are divided into three parts: (1) performance evaluation of six LMMs trained with different AltCLIP variants across two benchmarks, (2) top-1 positive example retrieval performance evaluation of six AltCLIP models, and (3) varying number of hard caption negatives.
-4.2.1 Setup
-We compare the following six types of AltCLIP models are as follows:
-  • Original AltCLIP: No additional data training.
-  • Random AltCLIP: Trained with randomly selected negative samples within a batch (batch size = 10) using 400K MAVIS data.
-  • Retrieved AltCLIP: Trained with 100K samples, each with 10 retrieval-based caption negatives.
-  • Rule-based AltCLIP: Trained with 100K samples, each with 10 rule-based caption negatives.
-  • Negative-images AltCLIP: Trained with 4K samples, each with 10 hard image negatives.
-  • Retrieved + Rule-based + Negative-images: Trained on all three datasets combined.
-The Random AltCLIP follows OpenAI’s training approach Radford et al.
-(2021), while the Retrieved, Rule-based, Negatives-images and Retrieved + Rule-based + Negatives-images AltCLIP models employ our proposed MMCLIP training approach.
-In the second and third stages, LMMs are trained on 67K G-LLaVA alignment data and 17K open-ended geometric reasoning problems.
-For the retrieval experiments, we assess the retrieval performance of trained AltCLIP models by selecting positive samples from a large pool of negative samples.
-Retrieval performance is evaluated using the Hit@1 metric.
-We construct four types of negative datasets: Random Negative: Caption negatives randomly selected in a training batch.
-Retrieval Negative: Caption negatives retrieved using the SimANS Zhou et al.
-(2022) model, selecting the top 100 most similar captions to the positive sample.
-Rule-based Negative: Caption negatives generated according to the rules defined in section 2.2.
-Image Negative: Image negatives generated from real-world geometric problems by MM-MATH.
-For the number of negative experiments, we investigate the impact of hard caption negative samples on MMGeoLM varying from 5 to 50.
-4.2.2 Results
-As shown in Figure 4, LMMs trained with hard negative samples consistently improve performance on the GeoQA and MM-MATH benchmarks.
-Notably, on MM-MATH, Negative-images AltCLIP, trained with only 4K image negative samples, surpasses Retrieved AltCLIP and Rule-based AltCLIP by 0.2% and 1.7%, respectively.
-On MM-MATH, Rule-based, Retrieved, and Image-negative AltCLIP consistently outperform Random AltCLIP on easy and medium-level questions.
-These results confirm the effectiveness of both hard negative construction methods.
-Furthermore, AltCLIP trained with a combination of caption and image negatives achieves the best performance on both GeoQA and MM-MATH, demonstrating their complementary effect in improving model performance.
-Further experiments on image negatives (Figure 4) reveal that Negative-image-MM-MATH-Align, constructed from real-world images, outperforms Negative-image-Mavis.
-This suggests that high-quality, realistic image negatives have a more pronounced impact on model performance than large-scale synthetic counterparts.
-The results of the retrieval performance experiments are shown in Figure 5.
-Under the negative-images setting, Negative-images AltCLIP achieves a retrieval score of 23%, while other AltCLIP variants remain below 15%.
-These results highlight the difficulty of retrieving realistic images.
-Notably, the dataset includes only 4,021 image negative samples, which is likely insufficient to fully exploit the model’s performance potential.
-In the retrieval-based and rule-based negative settings, Random AltCLIP outperforms Rule-based AltCLIP in retrieval-based settings and exceeds Retrieved AltCLIP in rule-based settings.
-These results indicate that while hard negative training enhances performance within specific domains, it may limit generalization.
-Furthermore, pre-training on random negatives followed by hard negatives improves both generalization and domain-specific performance, ultimately enhancing overall model effectiveness.
-Figure 6: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative samples.
+Compared to using the w/o MMCLIP (AltCLIP trained only with random negatives in a batch), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
+Ablation Study 1: Different Negatives Types
+Vision Encoder #Num. GeoQA $M$-MATH
+Easy Med Hard Avg
+Original - 45.4 40.5 20.8 4.5 23.8
+$\text{Random}_{\text{10}}$ 400K 45.0 39.8 22.7 4.5 24.9
+$\text{Retrieval}_{\text{10}}$ 100K 53.0 43.6 24.8 4.5 26.6
+$\text{Rule}_{\text{10}}$ 100K 48.8 45.4 27.6 4.5 28.1
+$\text{Image-based}_{\text{10}}$ 4K 54.9 45.6 26.5 4.5 29.0
+$\text{Retrieval}_{\text{10}}$$\text{Rule}_{\text{10}}$ 200K 56.5 47.5 28.3 4.5 29.4
+All Negatives 204K 58.2 49.8 29.1 4.5 30.4
+Table 3: Ablation study on vision encoder training with different types and quantities of hard negatives. The bottom panel reports performance when combining multiple types.
+To further evaluate the performance of vision encoders trained with different types of hard negatives, we compare seven negatives summarized in Table 3.
+All models are based on AltCLIP.
+The Original vision encoder is directly initialized without any additional training, while $\text{Random}_{\text{10}}$ follows the conventional in-batch random negatives sampling strategy (Radford et al.
+2021).
+The remaining five vision encoders are trained with our proposed MMCLIP method.
+The subscript 10 indicates a 1:10 positive-to-negative ratio.
+Results
+As shown in Table 3, vision encoders trained with hard negatives consistently outperform random in-batch negatives.
+even with only 4K image-based negatives , the $\text{Image-based}_{\text{10}}$ encoder achieves the highest average accuracy (29.0%) among all single hard negative methods on MM-Math.
+Furthermore, while $\text{Retrieval}_{10}$ and $\text{Rule}_{10}$ negatives individually provide moderate gains, their combination further improves performance (29.4%), indicating their complementary nature.
+The best results of All Negatives are achieved when all three types of negatives are combined, demonstrating that diverse hard negatives collectively enhance the model’s geometric understanding.
+These findings underscore the importance of both the type and diversity of hard negatives in contrastive training for vision encoders.
+Ablation Study 2: Image-based Negatives
+Method #Num. GeoQA $M$-Math
+Easy Med Hard Avg
+Rule-Img 100K 52.9 44.0 25.4 1.0 28.3
+Perturb-Img 4K 54.9 45.6 26.5 4.5 29.0
+$\Delta$ (Perturb-Rule) -96K +2.0 +1.6 +1.1 +3.5 +0.7
+Table 4: Effect of different image negative construction strategies. $\Delta$ rows give absolute gains of Perturb-Img over Rule-Img.
+2.1. Code Perturbation vs. Rule-design: Assessing Image Negatives Construction Methods
+We contrast two image–based negative construction methods.
+Perturb-Img applies our code perturbation method to 4K images, generating negative images that preserve similar structure while introducing subtle differences from the original diagrams.
+In contrast, Rule-Img is built from manually designed templates in the MAVIS corpus, enabling large-scale expansion.
+We construct 100K samples and retrieve most similar images as hard negatives.
+To ensure a fair comparison, we fix the positive-to-negative ratio at 1:10 in both settings.
+Results in Table 4 show that the 4K Perturb-Img setting exceeds the 100K Rule-Img on both GeoQA (54.9 vs.
+52.9) and MM-Math (29.0 vs.
+28.3).
+These results highlight that code-perturbed negatives, despite being smaller in scale, outperform synthetic data due to their semantic diversity and closer resemblance to real-world tasks.
+2.2. Length/Angle Marked Matters? Cross-Modal Consistency Analysis
+Figure 4: An example of vision encoder training data with or without numeric markings.
+Visual Marking GeoQA $M$-MATH
+Easy Med Hard Avg
+w/ length/angle/etc. 54.9 45.6 26.5 4.5 29.0
+w/o length/angle/etc. 51.7 50.0 25.1 4.5 28.6
+Table 5: Impact of numeric markings on vision encoder training. w/ and w/o denote settings with and without such numeric annotations, respectively.
+[…]
+2.3. Hard Negative Performance Gains or Trade-off for In-Batch Training
+Negative Strategy $M$-MATH
+Easy Med Hard Avg
+In-Batch 41.8 23.7 4.5 25.5
++image-based NS 51.7 (+9.9↑) 29.0 (+5.3↑) 4.5 32.1 (+6.6↑)
+Table 6: Comparison of In-Batch and In-Batch training with the proposed image-based negative strategy. Mean score over 5 random seeds on 4K image-based negatives.
+We leverage the proposed image-based negative strategy for the in-batch training to verify the effectiveness.
+In the in-batch negatives setting, each sample in the batch serves as a hard negative for others, with traditional in-batch typically selected at random.
+For a fair comparison, both methods use 10 negative examples per positive sample.
+As shown in Table 6, in-batch training with our proposed image-based negative strategy outperforms In-Batch on the easy, medium, and average subsets under the same amount of training data.
+This highlights the advantage of explicitly constructed image-based negatives in guiding the model to better distinguish visually similar examples, thus achieving higher overall accuracy.
+Furthermore, these results indicate that carefully selected mutual hard negatives within batches lead to more effective contrastive learning.
+Ablation Study 3: Training Strategy and Robustness
+Figure 5: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative ratio.
+3.1. Data Scaling: More Hard Negatives Are Not Always Better
+Our proposed MMGeoLM method allows flexible scaling of negatives.
+However, whether performance improves proportionally with an increasing number of hard negatives remains unclear.
+To address this, we investigate the impact of varying the number of retrieval-based negatives per sample from 5 to 50.
 As shown in Figure 5, MMGeoLM’s performance on MM-Math improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point.
-These results indicate that hard caption negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
-Training with only five hard negative samples outperforms training with ten randomly selected negatives (as shown in Table 1), demonstrating that hard negatives enhance multimodal mathematical reasoning more effectively.
-5 Related Work
+These results indicate that hard negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
+3.2. Robustness to Geometric Image Transformations
+Transformers Negatives $M$-MATH
+Easy Med Hard Avg
+No-Changing Text-based 48.6 29.1 9.0 31.9
+[…]
+Downscaling Text-based 45.6 26.8 0.0 28.4
+No-Changing Image-based 45.6 26.5 4.5 29.0
+[…]
+Downscaling Image-based 48.2 25.6 4.5 28.7
+Table 7: Robustness evaluation of MMGeoLM under image transformers (rotation and downscaling).
+Diagrams in real exams frequently undergo transformations, such as rotation and scaling.
+To evaluate MMGeoLM’s robustness to these variations, we apply two image perturbations: (1) a $30^{\circ}$ clockwise rotation, and (2) downscaling images by a factor of 0.5.
+Table 7 summarizes the results under both text-based and image-based negative conditions.
+Model performance remains largely stable under a $30^{\circ}$ rotation, showing only a minor accuracy drop of 0.9 (31.9 $\rightarrow$ 31.0) in the text-based setting.
+This suggests the model has adequate robustness against moderate image rotations.
+In contrast, downscaling significantly impacts accuracy, dropping text-based performance from 31.9 to 28.4.
+This decline indicates that textual captions are particularly sensitive to reduced image resolution whereas the image-based negative demonstrates stronger robustness, benefiting from diagram scale variations.
+Related Work
 Mathematics-related research has recently received significant attention in LLMs and LMMs.
-For text-only mathematical reasoning, several works employ external tools like Tora Gou et al.
-(2023) or intermediate step decomposition methods such as MAmmoTH Yue et al.
-(2023), Metamath Yu et al.
-(2023) and Math-Shepherd Yue et al.
-(2023).
-For multimodal mathematics tasks involving images, benchmarks have been proposed, including mathematical competition-oriented OlympiadBench He et al.
-(2024), geometry-focused datasets like Geometry3K Lu et al.
-(2021), VisScience Jiang et al.
-(2024), UniGeo Chen et al.
-(2022a) and GeoQA Chen et al.
-(2022a).
-Many works take effort on training multimodel math, including G-llava Gao et al.
-(2023), Meta-LLaVA Shi et al.
-(2024) and MAVIS Zhang et al.
-(2024c).
-However, these multimodal approaches typically suffer from inaccuracies in recognizing geometric elements, where our paper specifically addresses the enhancement of visual recognition accuracy.
-Vision encoders play a crucial role in understanding image content.
-While previous works Zhang et al.
-(2024a); Doveh et al.
-(2023a, b); Singh et al.
-(2023) have leveraged negative captions to enhance image comprehension, but not address the recognition of geometric elements within images.
-Methods such as NegCLIP Yuksekgonul et al.
-(2022) and TriCLIP Yang et al.
-(2024) focus on learning from negative samples, but their applicability is limited by the constrained number of negative samples.
-In this work, we propose a robust negative learning strategy that explicitly constructs negative samples for geometric elements while enabling arbitrary expansion, thereby improving geometric understanding in multimodal tasks.
-6 Conclusion
-This paper introduced MMCLIP, a novel hard negative contrastive learning framework to enhance LMMs’ geometric reasoning by refining their vision encoders with systematically generated image and text-based negatives.
+For text-only mathematical reasoning, several works employ external tools like Tora (Gou et al.
+2023) or intermediate step decomposition methods such as MAmmoTH (Yue et al.
+2023), Metamath (Yu et al.
+2023) and Math-Shepherd (Yue et al.
+2023).
+Many multimodal mathematics benchmarks have been proposed, including mathematical competition-oriented OlympiadBench (He et al.
+2024), geometry-focused datasets like Geometry3K (Lu et al.
+2021), VisScience (Jiang et al.
+2024), UniGeo (Chen et al.
+2022a) and GeoQA (Chen et al.
+2022a).
+Many works take effort on training multimodel math, including G-llava (Gao et al.
+2023), Meta-LLaVA (Shi et al.
+2024) and MAVIS (Zhang et al.
+2024b).
+However, these models often struggle with fine-grained recognition of geometric elements.
+Vision encoders are critical for capturing spatial and structural information in multimodal tasks.
+Prior works (Zhang, Awal, and Agrawal 2024; Doveh et al.
+2023a, b; Singh et al.
+2023) enhance image understanding using negative captions, but rarely target fine-grained geometric recognition.
+Methods such as NegCLIP (Yuksekgonul et al.
+2022) and TriCLIP (Yang et al.
+2024) focus on learning from negative samples, yet their effectiveness is limited by small or randomly selected negatives.
+In this work, we propose a scalable and geometry-aware negative learning strategy that explicitly constructs task-relevant hard negatives to improve geometric understanding.
+Conclusion
+This paper introduced two types of hard negatives, image-based and text-based negatives, targeted at geometric element understanding to enhance LMMs’ geometric reasoning.
 Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks.
-This work demonstrates the power of tailored hard negative sampling for developing LMMs with fine-grained visual understanding in specialized domains such as geometric understanding and reasoning.
-Limitations
-Despite the effectiveness of the proposed image hard negatives constructed by perturbing Python scripts, the method heavily relies on the accuracy of LLM-generated code and captions.
-Any systematic biases in these synthetic constructions may introduce artifacts that differ from human-designed geometric problems.
-As a result, the model’s performance on real-world visual inputs remains uncertain and requires further validation on diverse real-world datasets.
-Appendix A Prompts
-After multiple experimental iterations, we constructed two different types of prompts for our experiments, as illustrated in Figures 7 and 8.
-Specifically, Figure 7 demonstrates the prompt constructed by inputting the textual description and answer of realistic geometry problems to obtain the corresponding geometric diagram Python script.
-Meanwhile, Figure 8 illustrates the prompt obtained by modifying captions from positive examples to create slightly perturbed negative samples.
-Figure 7: Python script prompt from question and true answer in geometric problems
-Figure 8: Negative caption generation prompt from positive captions
-Appendix B Figure Comparison
-Figure 9 presents illustrative examples of image negative sample, covering analytical geometry and planar geometry.
-Geometric elements include triangles, quadrilaterals, as well as various relational properties such as perpendicularity and intersection.
-Additionally, certain metric properties absent from the original diagrams are explicitly indicated.
-Specifically, the first column shows the original diagrams, and the second column depicts diagrams generated from the given questions and the true answer, which, despite slight discrepancies, accurately capture the overall outlines of the original diagrams.
-The third and fourth columns display the constructed negative diagram instances—generated based upon their corresponding positive examples—demonstrating notable similarities but also clear and meaningful distinctions.
-Figure 9: Comparison of generated geometric figures.
-The first column shows the original images, the second column presents the positive samples, and the last two columns illustrate the constructed negative samples.
-Appendix C Case Study
-Figure 10: Comparison of GPT-4o and MMGeoLM in geometric problem-solving.
-MMGeoLM correctly answers the problem using a different solution compared to True Answer, while GPT-4o generates a non-existent element $\triangle ABD$.
-Figure 11: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
-C.1 Case Study
-In Figures 10 and 11, we compare the solutions generated by GPT-4o and our MMGeoLM.
-In Figure 10, the red text highlights GPT-4o’s misrecognition of two triangles in the image, leading to an incorrect final result.
-In contrast, MMGeoLM produces a correct solution, albeit through a reasoning process different from the ground-truth answer.
-This highlights MMGeoLM’s capability to generate diverse problem-solving approaches.
-Figure 11 illustrates errors in image-based reasoning for both MMGeoLM and GPT-4o.
-While both models make mistakes, MMGeoLM’s reasoning aligns more closely with the correct solution, demonstrating improved mathematical reasoning.
-Overall, our trained MMGeoLM model demonstrates greater reasoning diversity and enhanced geometric problem-solving performance compared to current mainstream models.
+This work demonstrates the power of hard negative contrastive learning for improving LMMs with fine-grained visual understanding.
+Extensive ablation studies confirm the effectiveness of hard negatives, highlighting the crucial role advancing in multimodal mathematical reasoning.

### v3

@@
-Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
+MMGeoLM: Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models
 Abstract
-Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks.
-However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving.
-To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity.
-We train CLIP using our hard negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving.
+Large Multimodal Models (LMMs) typically build on ViTs (e.g., CLIP), yet their training with simple random in-batch negatives limits the ability to capture fine-grained visual differences, particularly in geometric scenarios.
+To address this challenge, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity.
+We train a vision encoder (CLIP) using our hard negative training method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving.
 Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks.
@@
 We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations.
-These analyses yield important insights into optimizing hard negative strategies for geometric reasoning tasks.
+These analyses yield important insights into optimizing the training pipeline of vision encoder for fine-grained geometric reasoning tasks.
 https://github.com/THU-KEG/MMGeoLM.
-Introduction
-Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic 2024; OpenAI 2023; Bai et al.
-2023).
-It requires fine-grained identification of visual elements (Lu et al.
-2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al.
-2024).
-However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al.
-2024; Yang et al.
-2023; Goel et al.
-2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning.
+11footnotetext: Equal Contribution.
+1 Introduction
+Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic, 2024; OpenAI, 2023; Bai et al., 2023).
+It requires fine-grained identification of visual elements (Lu et al., 2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al., 2024).
+However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al., 2024; Yang et al., 2023; Goel et al., 2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning.
 Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers.
-As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI 2024a), Claude-3 (Anthropic 2024), and Qwen2.5-VL (Bai et al.
-2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
-[…]
-To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al.
-2023; Zhang et al.
-2024b; Peng et al.
-2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al.
-2024; Wang et al.
-2024) by aligning images with corresponding captions.
-However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al.
-2024a; Doveh et al.
-2023a).
+As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI, 2024a), Claude-3 (Anthropic, 2024), and Qwen2.5-VL (Bai et al., 2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
+To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al., 2023; Zhang et al., 2024c; Peng et al., 2024a; b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al., 2024; Wang et al., 2024) by aligning images with corresponding captions.
+However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al., 2024b; Doveh et al., 2023a).
 To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations.
 Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
+Figure 1: Examples of hallucination: non-existent elements and relation misinterpretation.
 Model Training Training Data $M$-Math
 AltCLIP No No 23.8
-AltCLIP In-Batch Randomly-sampled Negative (400K) 24.6
-AltCLIP MMCLIP ①: Retrieval-based Negative (100K) 26.6
-AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
-AltCLIP MMCLIP ③: Image-based Negative (4K) 28.2
+AltCLIP In-Batch Randomly-sampled Neg (400K) 24.6
+AltCLIP MMCLIP ①: Retrieval Neg (100K) 26.6
+AltCLIP MMCLIP ②: Rule Neg (100K) 28.1
+AltCLIP MMCLIP ③: Image Neg (4K) 28.2
 AltCLIP MMCLIP ①+② 28.4
@@
 Figure 2: Image-based and text-based hard negative construction and the corresponding MMCLIP training method.
-Method
-Negative Images Construction
+2 Hard Negative Data Construction
+2.1 Negative Images Construction
 We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives.
-Existing generation methods (Zhang et al.
-2024b; Zou et al.
-2024; Wei et al.
-2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems.
-In contrast, authentic geometry problems (Sun et al.
-2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock, Jiang, and July 2002).
-Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images.
+Existing generation methods (Zhang et al., 2024c; Zou et al., 2024; Wei et al., 2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems.
+In contrast, authentic geometry problems (Sun et al., 2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock et al., 2002).
+Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI, 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images.
 As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption.
 Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
-To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al.
-2024) to correct syntax issues when the code fails.
-Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind 2025).
-Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
+To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al., 2024) to correct syntax issues when the code fails.
+Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind, 2025).
+Upon evaluating 123 generated captions, three independent annotators confirmed that the accuracy of captions in representing geometric elements reaches 100%.
+Further details are provided in Appendix F.
 To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically.
@@
 These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
-Negative Captions Construction
+Representative examples of the constructed image-based negatives are provided in Appendix B.
+2.2 Negative Captions Construction
 We propose two methods for constructing text-based hard negative samples.
@@
 Retrieval-based Hard Negative.
-Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al.
-2020; Xiong et al.
-2020; Huang et al.
-2020).
-However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al.
-2020; Zhou et al.
-2022; Yang et al.
-2024).
-To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al.
-2024b), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption.
-We use SimANS (Zhou et al.
-2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
+Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al., 2020; Xiong et al., 2020; Huang et al., 2020).
+However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al., 2020; Zhou et al., 2022; Yang et al., 2024).
+To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al., 2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption.
+We use SimANS (Zhou et al., 2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
 Rule-based Hard Negative.
-Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al.
-2022; Zhang, Awal, and Agrawal 2024), while others leverage in-context learning to generate meaningful modifications (Patel et al.
-2025) .
+Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al., 2022; Zhang et al., 2024a), while others leverage in-context learning to generate meaningful modifications (Patel et al., 2025).
 However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
-To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al. 2024) and identify four major types of image element recognition errors:
+To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al., 2024) and identify four major types of image element recognition errors:
   1. 1.
@@
     Numerical values: Adjusting numerical values in captions, such as modifying angles or segment lengths.
-To address these error types, we design rule-based strategies that use GLM-4 (GLM et al. 2024) to modify each positive caption and generate 10 hard negative examples.
+To address these error types, we design rule-based strategies that use GLM-4 (GLM et al., 2024) to modify each positive caption and generate 10 hard negative examples.
 These text-based negatives are used in model training, and two representative types are shown in Figure 2.
-Hard Negative CLIP Training
+3 MMGeoLM: Architecture and Training
 Figure 3: Overview of the MMGeoLM training pipeline, including main and each ablation experiment configurations, training strategies, and datasets.
-Though CLIP-trained vision encoders are widely used in LMMs (Radford et al.
-2021; Doveh et al.
-2023b), they typically combine random examples for a batch, which limits their capacity for fine-grained image understanding during training.
-While methods like NegCLIP (Yuksekgonul et al.
-2022) introduce hard negatives, they still rely heavily on randomly selected negatives within a batch.
-Considering the methods in the above section have allowed us to construct a large number of hard negatives, we propose a MMCLIP training strategy, which focuses on each single sample and only uses its corresponding hard negatives for training.
-In this way, the vision encoder is forced to distinguish subtle differences, thereby enhancing its fine-grained geometry understanding capacities.
-Formally, given a text-based negative training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive samples and $T_{j}^{-}$ is the negative samples conducted around the image.
-Our loss function is formulated as follows:
-$\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
+3.1 Architecture
+We adopt the LLaVA architecture (Liu et al., 2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone.
+For the LLM backbone, we use MAmmoTH2-7B (Yue et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025).
+The vision encoder is based on AltCLIP (Chen et al., 2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
+3.2 Training Pipeline
+[…]
+3.3 Hard Negative CLIP Training
+In-batch Training.
+Conventional CLIP-trained vision encoders (Radford et al., 2021; Doveh et al., 2023b) typically combine random examples for a batch and adopt in-batch training loss over all the samples in each batch.
+For each batch containing samples $\{(I_{i},T_{i})_{i=1}^{N}\}$, where each $I_{i},T_{i}$ represents a pair of image and text caption, the loss is calculated as
+$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left[-\ln\frac{\exp(s(I_{i},T_{i}))}{\sum_{j}\exp(s(I_{i},T_{j}))}-\ln\frac{\exp(s(I_{i},T_{i}))}{\sum_{j}\exp(s(I_{j},T_{i}))}\right],$ (1)
 where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings.
-Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ is expected to be higher than a randomly chosen one.
+MMCLIP Training.
+In our data construction process, hard negative captions and images are generated for each image-caption pair without providing corresponding positive samples for these hard negatives.
+As a result, our MMCLIP training derives the loss solely from one positive sample and its associated hard negatives within each batch.
+This guides the model to discern subtle differences, enhancing its fine-grained geometric understanding.
+Formally, given a text-based negative training batch $\{I,T^{+},T_{i}^{-}|_{i=1}^{N}\}$, where $I$ represents the image, $T^{+}$ denotes the positive caption and $T_{i}^{-}$ is the negative caption conducted around the image.
+The loss function is formulated as follows:
+$\mathcal{L}=-\ln\frac{\exp(s(I,T^{+}))}{\exp(s(I,T^{+}))+\sum_{i=1}^{N}\exp(s(I,T_{i}^{-}))}.$ (2)
+Since negative samples are constructed around elements within the image, their score $s(I,T_{i}^{-})$ is expected to be higher than a randomly chosen one.
 From the above equation, we calculate the gradient for positive and negative samples:
-$\frac{\partial\mathcal{L}}{\partial s_{i}^{+}}=-\frac{\sum_{j=1}^{N}\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (2)
-$\frac{\partial\mathcal{L}}{\partial s^{-}_{j}}=\frac{\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (3)
-where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$.
-Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment.
-The training strategy is illustrated in Figure 2.
-Unlike in-batch training, which relies on batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for rule-based or retrieval-based unimodal negatives.
+$\frac{\partial\mathcal{L}}{\partial s^{+}}$ $=-\frac{\sum_{i=1}^{N}\exp(s^{-}_{i})}{\exp(s^{+})+\sum_{i=1}^{N}\exp(s^{-}_{i})}$
+$\frac{\partial\mathcal{L}}{\partial s^{-}_{i}}$ $=\frac{\exp(s^{-}_{i})}{\exp(s^{+})+\sum_{i=1}^{N}\exp(s^{-}_{i})}$
+where $s^{+}=s(I,T^{+})$ and $s_{i}^{-}=s(I,T_{i}^{-})$.
+Since the number of hard negatives is unrestricted, the gradient functions can be effectively optimized, leading to improved image-caption alignment.
+The training strategy is illustrated in Figure 2, where we also provide image-based negatives training batches $\{T,I^{+},I_{i}^{-}|_{i=1}^{M}\}$, the loss can be derived similarly to Equation 2.
+Unlike in-batch training, which relies on random batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for unimodal negatives, such as rule-based or retrieval-based.
 We further validate different training strategies only for hard negatives in the ablation experiments.
-Supervised Fine-tuning Data
-[…]
-Multimodal Mathematical Training
-Architecture
-We adopt the LLaVA architecture (Liu et al.
-2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone.
-For the LLM backbone, we use MAmmoTH2-7B (Yue et al.
-2024) and Qwen2.5-7B-Instruct (Qwen et al.
-2025).
-The vision encoder is based on AltCLIP (Chen et al.
-2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
-Training Details
+4 Experiments
+To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs.
+Additionally, we analyze the impact of different vision encoder types, image-based negatives and training strategies in the subsequent ablation studies.
+4.1 Main Experiment: MMGeoLM Evaluation
+4.1.1 Datasets and Metrics
+We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended.
+The multiple-choice benchmark includes GeoQA (Chen et al., 2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al., 2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al., 2023), widely used for evaluating LMMs’ performance.
+The open-ended benchmark is MM-Math (Sun et al., 2024), which features high discriminative difficulty sourced from secondary school-level problems.
+We use accuracy (ACC) as the metric.
+For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems.
+For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
+4.1.2 Baselines
+The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic, 2024), Claude-3.5-Sonnet, GPT-4o-20240513 (OpenAI, 2024a), and GPT-4V.
+Open-source LMMs: MAVIS-7B (Zhang et al., 2024c), G-LLaVA-7B (Gao et al., 2023), InternVL2-8B (Chen et al., 2024), LLaVA-OneVision-7B (Li et al., 2024), Qwen2.5-VL-7B-Instruct (Bai et al., 2025), Chimera-Reasoner-8B (Peng et al., 2024b), InternLM-XComposer2-7B (Dong et al., 2024), Phi-3-Vision-128K-Instruct (Abdin et al., 2024), and Math-LLaVA-13B (Shi et al., 2024).
+For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al., 2021; Lu et al., 2023).
+Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al., 2024c; Sun et al., 2024).
+4.1.3 Overall Results
 Model GeoQA MathVista We-Math $M$-Math
-GPS ALG S2 S3 Easy Med Hard Avg
+GEO ALG S2 S3 Easy Med Hard Avg
 Human* 92.3 48.4 50.9 - - 90.7 81.9 47.6 80.4
@@
 Table 2: Accuracy (%) of various LMMs on GeoQA, MathVista, We-Math, and $M$-Math. ‘S2/S3’ denotes the two-step settings in We-Math. Results marked * are taken from their original papers.
-Our training pipeline, illustrated in Figure 3, consists of three stages covering both main and ablation experiments.
-In the first stage, we perform vision-text alignment training.
-For text-based hard negatives, we use the Mavis image-caption alignment dataset (Zhang et al.
-2024b).
-For image-based negatives, we collect 4K geometry questions from real middle-school exams.
-Additional details, including the types of negatives and the negative-to-positive ratio, are provided in Figure 3.
-In the second stage, we tune the MLP adapter using 67K image-text pairs from the G-LLaVA dataset (Gao et al.
-2023).
-In the third stage, we conduct supervised fine-tuning.
-For main experiments, we use a combined dataset comprising 300K Mavis instruction data (Zhang et al.
-2024b), 117K G-LLaVA instruction data (Gao et al.
-2023), and 17K open-ended geometry questions.
-For ablation studies, we only use the 17K newly collected geometry problems for efficient validation.
-We name the final trained model as MMGeoLM.
-Experiments
-To quantitatively evaluate the effectiveness of our MMGeoLM in geometric problem-solving, we conduct comparative experiments on four geometric benchmarks using various LMMs.
-Additionally, we analyze the impact of different vision encoder types, image-based negatives and training strategies in the subsequent ablation studies.
-LMM Geometric Reasoning Evaluation
-Datasets and Metrics
-We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended.
-The multiple-choice benchmark includes GeoQA (Chen et al.
-2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al.
-2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al.
-2023), widely used for evaluating LMMs’ performance.
-The open-ended benchmark is MM-Math (Sun et al.
-2024), which features high discriminative difficulty sourced from secondary school-level problems.
-We use accuracy (ACC) as the metric.
-For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems.
-For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
-Baselines
-The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic 2024), Claude-3.5-Sonnet, GPT-4o (OpenAI 2024a), and GPT-4V.
-Open-source LMMs: MAVIS-7B (Zhang et al.
-2024b), G-LLaVA-7B (Gao et al.
-2023), InternVL2-8B (Chen et al.
-2024), LLaVA-OneVision-7B (Li et al.
-2024), Qwen2.5-VL-7B-Instruct (Bai et al.
-2025), Chimera-Reasoner-8B (Peng et al.
-2024b), InternLM-XComposer2-7B (Dong et al.
-2024), Phi-3-Vision-128K-Instruct (Abdin et al.
-2024), and Math-LLaVA-13B (Shi et al.
-2024).
-For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al.
-2021; Lu et al.
-2023).
-Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al.
-2024b; Sun et al.
-2024).
-Overall Results
 As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks.
 On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B.
-As Chimera-Reasoner-8B was trained on GeoQA (Peng et al.
-2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset.
+As Chimera-Reasoner-8B was trained on GeoQA (Peng et al., 2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset.
 The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach.
@@
 In contrast, hard problems involve multi-step reasoning, where its impact is more limited.
-Compared to using the w/o MMCLIP (AltCLIP trained only with random negatives in a batch), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
-Ablation Study 1: Different Negatives Types
+Compared to using the w/o MMCLIP (In-batch trained AltCLIP with random negatives), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
+To rule out the possibility that performance gains stem from training–test leakage, we perform a contamination analysis (Appendix E) and we confirm that the evaluation sets contain no near-duplicate items.
+4.2 Ablations
+4.2.1 Ablation Study I: Different Negative Types
 Vision Encoder #Num. GeoQA $M$-MATH
@@
 All models are based on AltCLIP.
-The Original vision encoder is directly initialized without any additional training, while $\text{Random}_{\text{10}}$ follows the conventional in-batch random negatives sampling strategy (Radford et al.
-2021).
-The remaining five vision encoders are trained with our proposed MMCLIP method.
-The subscript 10 indicates a 1:10 positive-to-negative ratio.
-Results
+The Original vision encoder is directly initialized without any additional training, while $\text{Random}_{\text{10}}$ follows the conventional in-batch random negatives sampling strategy in Section 3.3.
+The remaining five vision encoders are trained with our proposed MMCLIP method in Section 3.3.
+The subscript ‘10’ indicates a 1:10 positive-to-negative ratio.
+We further complement this analysis with retrieval evaluations reported in Appendix C, using the Hit@1 metric on four constructed evaluation sets.
+Results.
 As shown in Table 3, vision encoders trained with hard negatives consistently outperform random in-batch negatives.
@@
 These findings underscore the importance of both the type and diversity of hard negatives in contrastive training for vision encoders.
-Ablation Study 2: Image-based Negatives
-Method #Num. GeoQA $M$-Math
+4.2.2 Ablation Study II: Image-based Negatives
+Method #Num GeoQA $M$-Math
 Easy Med Hard Avg
-Rule-Img 100K 52.9 44.0 25.4 1.0 28.3
-Perturb-Img 4K 54.9 45.6 26.5 4.5 29.0
-$\Delta$ (Perturb-Rule) -96K +2.0 +1.6 +1.1 +3.5 +0.7
-Table 4: Effect of different image negative construction strategies. $\Delta$ rows give absolute gains of Perturb-Img over Rule-Img.
-2.1. Code Perturbation vs. Rule-design: Assessing Image Negatives Construction Methods
-We contrast two image–based negative construction methods.
-Perturb-Img applies our code perturbation method to 4K images, generating negative images that preserve similar structure while introducing subtle differences from the original diagrams.
-In contrast, Rule-Img is built from manually designed templates in the MAVIS corpus, enabling large-scale expansion.
-We construct 100K samples and retrieve most similar images as hard negatives.
-To ensure a fair comparison, we fix the positive-to-negative ratio at 1:10 in both settings.
-Results in Table 4 show that the 4K Perturb-Img setting exceeds the 100K Rule-Img on both GeoQA (54.9 vs.
-52.9) and MM-Math (29.0 vs.
-28.3).
-These results highlight that code-perturbed negatives, despite being smaller in scale, outperform synthetic data due to their semantic diversity and closer resemblance to real-world tasks.
-2.2. Length/Angle Marked Matters? Cross-Modal Consistency Analysis
-Figure 4: An example of vision encoder training data with or without numeric markings.
+$\text{Rule}_{img}$ 100K 52.9 44.0 25.4 1.0 28.3
+Perturb 4K 54.9 45.6 26.5 4.5 29.0
+$\Delta$ -96K +2.0 +1.6 +1.1 +3.5 +0.7
+Table 4: Effect of different image negative construction strategies. $\Delta$ rows give absolute gains of Perturb over $\text{Rule}_{img}$.
 Visual Marking GeoQA $M$-MATH
@@
 Table 5: Impact of numeric markings on vision encoder training. w/ and w/o denote settings with and without such numeric annotations, respectively.
+Code Perturbation vs.
+Rule-design: Assessing Image Negatives Construction Methods.
+We contrast two image–based negative construction methods.
+Perturb applies our code perturbation method to 4K images, generating negative images that preserve similar structure while introducing subtle differences.
+In contrast, $\text{Rule}_{img}$ is built from manually designed templates in the Mavis corpus.
+We construct 100K samples and retrieve the most similar images as hard negatives.
+To ensure a fair comparison, we fix the positive-to-negative ratio at 1:10 in both settings.
+Results in Table 5 show that the 4K Perturb setting exceeds the 100K $\text{Rule}_{img}$ on both GeoQA (54.9 vs.
+52.9) and MM‑Math (29.0 vs.
+28.3).
+These results highlight that code-perturbed negatives, despite being smaller in scale, outperform synthetic data due to their semantic diversity and closer resemblance to real-world tasks.
+This advantage is corroborated by additional analyses in Appendix D, where Perturb-based negatives are shown to have a more distinct distribution than $\text{Rule}_{img}$ counterparts and higher semantic similarity to real-world supervised data.
+Figure 4: An example of vision encoder training data with or without numeric markings.
+In-Batch Negative GeoQA $M$-MATH
+Easy Med Hard Avg
+Random 46.2 41.8 23.7 4.5 25.5
+Image-based NS 57.3 51.7 29.0 4.5 32.1
+Table 6: Comparison of in-batch training with random negatives vs. in-batch training with image-based negatives. Mean score over 5 random seeds on 4K image-based negatives.
 […]
-2.3. Hard Negative Performance Gains or Trade-off for In-Batch Training
-Negative Strategy $M$-MATH
-Easy Med Hard Avg
-In-Batch 41.8 23.7 4.5 25.5
-+image-based NS 51.7 (+9.9↑) 29.0 (+5.3↑) 4.5 32.1 (+6.6↑)
-Table 6: Comparison of In-Batch and In-Batch training with the proposed image-based negative strategy. Mean score over 5 random seeds on 4K image-based negatives.
+Hard Negative Performance Gains or Trade-off for In-Batch Training.
 We leverage the proposed image-based negative strategy for the in-batch training to verify the effectiveness.
@@
 Furthermore, these results indicate that carefully selected mutual hard negatives within batches lead to more effective contrastive learning.
-Ablation Study 3: Training Strategy and Robustness
-Figure 5: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative ratio.
-3.1. Data Scaling: More Hard Negatives Are Not Always Better
-Our proposed MMGeoLM method allows flexible scaling of negatives.
-However, whether performance improves proportionally with an increasing number of hard negatives remains unclear.
-To address this, we investigate the impact of varying the number of retrieval-based negatives per sample from 5 to 50.
-As shown in Figure 5, MMGeoLM’s performance on MM-Math improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point.
-These results indicate that hard negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
-3.2. Robustness to Geometric Image Transformations
+4.2.3 Ablation Study III: Training Strategy and Robustness
+Figure 5: MMGeoLM performance with varying numbers of hard negative ratio.
 Transformers Negatives $M$-MATH
@@
 Downscaling Image-based 48.2 25.6 4.5 28.7
-Table 7: Robustness evaluation of MMGeoLM under image transformers (rotation and downscaling).
-Diagrams in real exams frequently undergo transformations, such as rotation and scaling.
-To evaluate MMGeoLM’s robustness to these variations, we apply two image perturbations: (1) a $30^{\circ}$ clockwise rotation, and (2) downscaling images by a factor of 0.5.
-Table 7 summarizes the results under both text-based and image-based negative conditions.
-Model performance remains largely stable under a $30^{\circ}$ rotation, showing only a minor accuracy drop of 0.9 (31.9 $\rightarrow$ 31.0) in the text-based setting.
-This suggests the model has adequate robustness against moderate image rotations.
-In contrast, downscaling significantly impacts accuracy, dropping text-based performance from 31.9 to 28.4.
+Table 7: Robustness evaluation of MMGeoLM trained on text/image negatives under image transformers (rotation and downscaling).
+Data Scaling: More Hard Negatives Are Not Always Better.
+Our proposed MMGeoLM method allows flexible scaling of negatives.
+However, whether performance improves proportionally with an increasing number of hard negatives remains unclear.
+To address this, we investigate the impact of varying the number of retrieval-based negatives per sample from 5 to 50.
+As shown in Figure 5, MMGeoLM’s performance on MM-Math improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point.
+These results indicate that hard negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
+Robustness to Geometric Image Transformations.
+Diagrams in real exams often undergo transformations such as rotation or scaling.
+To evaluate MMGeoLM’s robustness, we apply two image perturbations: (1) a $30^{\circ}$ clockwise rotation and (2) downscaling by 0.5.
+Table 7 reports the results under both text-based and image-based negative conditions.
+Model performance remains largely stable under a $30^{\circ}$ rotation, showing only a minor accuracy drop of 0.9% in the text-based setting.
+This suggests the model is robust against moderate image rotations.
+In contrast, downscaling significantly impacts accuracy, dropping text-based performance from 31.9% to 28.4%.
 This decline indicates that textual captions are particularly sensitive to reduced image resolution whereas the image-based negative demonstrates stronger robustness, benefiting from diagram scale variations.
-Related Work
-Mathematics-related research has recently received significant attention in LLMs and LMMs.
-For text-only mathematical reasoning, several works employ external tools like Tora (Gou et al.
-2023) or intermediate step decomposition methods such as MAmmoTH (Yue et al.
-2023), Metamath (Yu et al.
-2023) and Math-Shepherd (Yue et al.
-2023).
-Many multimodal mathematics benchmarks have been proposed, including mathematical competition-oriented OlympiadBench (He et al.
-2024), geometry-focused datasets like Geometry3K (Lu et al.
-2021), VisScience (Jiang et al.
-2024), UniGeo (Chen et al.
-2022a) and GeoQA (Chen et al.
-2022a).
-Many works take effort on training multimodel math, including G-llava (Gao et al.
-2023), Meta-LLaVA (Shi et al.
-2024) and MAVIS (Zhang et al.
-2024b).
-However, these models often struggle with fine-grained recognition of geometric elements.
+5 Related Work
+Mathematics-related research has recently received significant attention in large models.
+For text-only mathematical reasoning, several works employ external tools like Tora (Gou et al., 2023) or intermediate step decomposition methods such as MAmmoTH (Yue et al., 2023), Metamath (Yu et al., 2023) and Math-Shepherd (Yue et al., 2023).
+Many multimodal mathematics benchmarks have been proposed, including mathematical competition-oriented OlympiadBench (He et al., 2024), geometry-focused datasets like Geometry3K (Lu et al., 2021), VisScience (Jiang et al., 2024), UniGeo (Chen et al., 2022a) and GeoQA (Chen et al., 2022a).
+Many works take effort on training multimodal math, including G-llava (Gao et al., 2023), Meta-LLaVA (Shi et al., 2024) and MAVIS (Zhang et al., 2024c).
+These models often struggle with fine-grained recognition of geometric elements.
 Vision encoders are critical for capturing spatial and structural information in multimodal tasks.
-Prior works (Zhang, Awal, and Agrawal 2024; Doveh et al.
-2023a, b; Singh et al.
-2023) enhance image understanding using negative captions, but rarely target fine-grained geometric recognition.
-Methods such as NegCLIP (Yuksekgonul et al.
-2022) and TriCLIP (Yang et al.
-2024) focus on learning from negative samples, yet their effectiveness is limited by small or randomly selected negatives.
+Prior works (Zhang et al., 2024a; Doveh et al., 2023a; b; Singh et al., 2023) enhance image understanding using negative captions, but rarely target fine-grained geometric recognition.
+Methods such as NegCLIP (Yuksekgonul et al., 2022) and TriCLIP (Yang et al., 2024) focus on learning from negative samples, yet their effectiveness is limited by small or randomly selected negatives.
 In this work, we propose a scalable and geometry-aware negative learning strategy that explicitly constructs task-relevant hard negatives to improve geometric understanding.
-Conclusion
-This paper introduced two types of hard negatives, image-based and text-based negatives, targeted at geometric element understanding to enhance LMMs’ geometric reasoning.
+6 Conclusion
+This paper introduces two types of hard negatives, image-based and text-based negatives, targeted at geometric element understanding to enhance LMMs’ geometric reasoning.
 Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks.
-This work demonstrates the power of hard negative contrastive learning for improving LMMs with fine-grained visual understanding.
-Extensive ablation studies confirm the effectiveness of hard negatives, highlighting the crucial role advancing in multimodal mathematical reasoning.
+Our ablation analyses corroborate three key findings.
+First, both image-based and text-based negatives enhance geometric problem-solving, and their effects are complementary when used jointly.
+Second, the proposed image-based negatives, constructed via code perturbation, exhibit advantages in terms of data efficiency, annotation robustness, and enhanced performance combined with traditional in-batch training methods.
+Third, increasing the number of negatives does not monotonically improve performance; instead, our MMCLIP training strategy demonstrates robustness by balancing the quality and diversity of negatives, yielding consistent gains across benchmarks.
+Ethics Statement
+Our 4K image-based negatives and 17K supervised fine-tuning training data were collected from the 21st Century Education. We have obtained an official authorization agreement from them for research.
+Reproducibility Statement
+The training dataset, 4K image-based negatives and 17K supervised fine-tuning training data, will be publicly released after the paper publication, with a download link to be provided in the camera-ready version.
+For the reproducibility of our training strategy, we provide three components.
+First, the implementation of MMCLIP training with both image-based and text-based negatives, along with the corresponding hyperparameters, is included in the supplementary material.
+Second, the MLP training setup follows the official implementation released in the LLaVA GitHub repository.
+Third, for incorporating LLMs, we adopt the training code and parameters from the official LLaVA GitHub repository for MAmmoTH2-7B, and we use the official code and configurations provided in the LLaVA-Next GitHub repository for Qwen2.5-7B.
+The Use of Large Language Models
+In this work, large language models were employed in three ways.
+First, they were used to generate code-based geometric diagrams and the corresponding captions for constructing image-based negatives.
+Second, they were applied to modify positive captions in order to derive negative captions for text-based negatives.
+Third, they were utilized to assist in polishing the writing of the manuscript.
+Appendix A Hard Negative Statistics
+We conduct a detailed analysis of the constructed hard negative datasets, covering the data sources, construction methods, the number of positive samples, the number of hard negatives per positive sample, and additional notes on the perturbation strategies.
+The statistics are summarized in Table 8.
+Data Type Source & Method #Positives #Hard Negatives Notes
+Retrieval‑based Negative MAVIS + SimANS 100K 10-50 Dense retrieval, high semantic similarity
+Rule‑based Negative MAVIS + GLM‑4 100K 10 Shape / relation / value perturbations
+Image‑based Negative New Collection + Gemini2.5 4K 10 Code‑driven geometric modifications
+Table 8: Statistics of constructed hard‑negative samples. “#Hard Negatives” indicates the number of negatives per positive sample.
+Appendix B Examples of Image-based Negatives
+Figure 6 presents some examples of image-based negatives, covering analytical geometry and planar geometry.
+Geometric elements include triangles, quadrilaterals, and various relational properties such as perpendicularity and intersection.
+Additionally, certain length/angle properties absent from the original diagrams are explicitly indicated.
+Specifically, the first column shows the original geometric images, and the second column depicts images generated from the given questions and the true answer, which, despite slight discrepancies, accurately capture the overall outlines of the original diagrams.
+The third and fourth columns display the constructed hard negative image based upon their corresponding negative captions and corresponding Python script—demonstrating notable similarities but clear and meaningful distinctions.
+Figure 6: Comparison of generated geometric figures.
+The first column shows the original images, the second column presents the positive images, and the last two columns illustrate the constructed hard negatives.
+Appendix C Retrieval Experiments for Vision Encoder
+In addition to ablation evaluation, retrieval performance offers a direct measure of image–text alignment, computed via similarity scores between paired inputs.
+As a result, the ranking position of the ground-truth positive sample reflects the quality of learned representations.
+To assess retrieval capabilities, we evaluate the trained AltCLIP models using the Hit@1 metric, which measures whether the correct caption is ranked highest among a large pool of candidates.
+Specifically, we construct four evaluation sets of hard negatives, each containing 500 samples, corresponding to the following strategies:
+  • Random Negative: Caption negatives randomly selected.
+  • Retrieval Negative: Caption negatives retrieved using the SimANS model, selecting the top 100 most similar captions to the positive sample.
+  • Rule-based Negative: Caption negatives generated according to the rules defined in earlier section.
+  • Image-based Negative: Image negatives generated from real-world geometric problems from new collection.
+Table 9 reports retrieval performance of various vision encoders trained with different types of negative samples in this experiment.
+Figure 7: Retrieval performance of AltCLIP variants trained with different strategies.
+AltCLIP Variant #Negative Source (K)
+Neg
+pos
+Train
+Rnd Ret Rule Img
+AltCLIP-Original
+AltCLIP-Rand 400 10 IB
+AltCLIP-Retrieval 100 10 $M$
+AltCLIP-Rule 100 10 $M$
+AltCLIP-Image 4 10 $M$
+AltCLIP-Rand+Ret 400 100 10 IB+$M$
+AltCLIP-Rand+Rule 400 100 10 IB+$M$
+AltCLIP-Rand+Img 400 4 10 IB+$M$
+AltCLIP-Ret+Rule 100 100 10 $M$
+AltCLIP-Ret+Rule+Img 100 100 4 10 $M$
+AltCLIP-All 400 100 100 4 10 IB+$M$
+Table 9: Dataset composition for each AltCLIP vision encoder variant.
+Columns list the number of positive pairs (in thousands) drawn from four negative-source types: Rnd = random, Ret = retrieval, Rule = rule-based, Img = image-based.
+Neg/pos indicates negatives-per-positive; training strategies: IB = In-batch, $M$ = MMCLIP, IB+$M$ = hybrid.
+The retrieval performance of different AltCLIP variants is summarized in Figure 7.
+Under the image-based negative setting, AltCLIP-Image achieves the highest Hit@1 score (23%), substantially outperforming all other variants, most of which remain below 15%.
+This demonstrates the superior alignment learned from visually grounded negatives.
+However, the performance is still constrained, likely due to the limited size of the image-based dataset (4K samples), which may restrict the model’s capacity to generalize under this setting.
+In retrieval- and rule-based evaluations, AltCLIP-Rand achieves the highest scores among single-source variants (64% and 31%, respectively), outperforming both AltCLIP-Retrieval and AltCLIP-Rule.
+This suggests that while domain-specific negatives help the model specialize, pre-training on random negatives provides broader generalization capabilities.
+Combining random negatives with domain-specific ones further improves robustness.
+For instance, AltCLIP-Rand+Rule yields strong performance across retrieval (41%) and rule-based (80%) settings, while AltCLIP-Rand+Ret+Rule+Img consistently achieves top scores across all four settings.
+These results indicate that starting with random negatives and progressively incorporating hard negatives can effectively balance generalization and domain-specific precision in image–text alignment.
+Appendix D Why Fewer Shot Image-Based Negatives Work: A Similarity and Distribution Study
+In above ablation study, we observe that only 4K image-based negatives generated can match or even surpass the performance of 100K retrieval-based negative image from MAVIS.
+To better understand this phenomenon, we conduct two follow-up analyses.
+Negative Type Max Cosine Similarity
+>0.90>0.90 >0.85>0.85 >0.70>0.70 >0.60>0.60
+Image‑based vs. SFT 20.89% 78.68% 98.43% 99.43%
+Retrieval‑based vs. SFT 0.00% 26.18% 98.00% 99.28%
+Table 10: Semantic similarity between hard negative captions and the 17K SFT dataset.
+Values denote the percentage of negative captions whose maximum cosine similarity with any SFT question exceeds the specified threshold.
+First, we evaluate the semantic similarity between negative samples and the supervised fine-tuning dataset.
+Specifically, we employ the sup-simcse-roberta-large model to encode (1) the captions of image-based negatives, (2) the top-10 retrieval-based negative captions, and (3) the questions from the 17K SFT dataset.
+We then compute the cosine similarity between each caption and the questions in SFT, and for each caption, we retain the maximum similarity score as its semantic alignment score.
+Next, we calculate the proportion of captions whose maximum similarity exceeds various thresholds (0.90, 0.85, 0.70, and 0.60).
+As shown in Table 10, image-based negatives exhibit substantially higher similarity to the SFT dataset, with 20.89% exceeding 0.90 and 78.68% exceeding 0.85.
+In contrast, retrieval-based negatives show much lower proportions, with only 0.00% and 26.18% exceeding these respective thresholds.
+These results suggest that image-based negatives, which are derived from real-world problems, exhibit closer semantic alignment with downstream supervised data compared to artificially constructed samples.
+Next, we visualized the embeddings of both image-based negatives and retrieval-based negatives using t-SNE, reducing them to two dimensions, as shown in Figure 8.
+The visualization clearly shows a distinct separation between the two types of negative samples.
+We further evaluate the separation of the two negative types using K-means clustering on the embeddings of both image-based and retrieval-based negatives.
+We then calculated two separation metrics: K-means Separation Accuracy and Natural Separation Score.
+The results are as Table 11:
+Figure 8: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
+Metric Value
+K-means Separation Accuracy 0.90
+Natural Separation Score 0.90
+Table 11: Clustering-based separation metrics between image-based and retrieval negative samples. Higher values indicate stronger distinguishability between the two groups in embedding space.
+From Table 11, we observe that both the K-means Separation Accuracy and the Natural Separation Score reach 0.90, indicating a strong degree of separability between the two types of negatives.
+This suggests that the 4K image-based negatives are not only distinguishable from retrieval-based ones, but also semantically closer to downstream tasks.
+Consequently, their effectiveness stems from their higher alignment with real-world data, rather than relying on large-scale rule-based construction.
+In contrast, datasets built purely through human-making rules may require a substantially larger volume to achieve comparable performance, resulting in lower data efficiency.
+Appendix E Training-Test Overlap and Data-Contamination Analysis
+Both the training and test datasets are drawn from genuine exam questions in an open‑ended format; thus, accidental duplicates could harm evaluation scores.
+To assess this risk, we embed every question with sup-simcse‑roberta‑large and compute the cosine similarity between each test question and the entire training datasets.
+For each test question we record its maximum and top‑5 similarity scores.
+Figure 9: Examples of top-5 most similar training–test question pairs used for contamination analysis.
+Each pair includes the test question, its top training matches, and the associated cosine similarity.
+Figure 9 shows the examples of top 5 similar score pair .
+For the top 1 score (sim=0.9977=0.9977), although their wording is close, a key numerical change ( $AB=5,DC=11$ vs.
+$AB=6,DC=13$) yields a absolute different solution path.
+To ensure a clean separation between training and test sets, we remove from the training data any question whose similarity with any test question exceeds $0.995$.
+Table 12 summarizes the resulting overlap statistics.
+Under this filtering strategy, 100% of the test questions have no near-duplicate counterparts in the training data, and over 97% have a maximum similarity below $0.990$, demonstrating that our evaluation is not compromised by training leakage.
+Similarity Threshold Percentage
+<0.995<0.995 100%
+<0.990<0.990 97.28%
+<0.980<0.980 90.56%
+<0.970<0.970 81.26%
+Table 12: Proportion of test questions whose maximum similarity to the training set falls below various thresholds, after filtering.
+Appendix F Human Verification of Generated Captions
+We randomly sampled 123 captions generated by the LLM call from the pipeline of image-based negatives.
+The sample size was chosen to cover diverse geometric configurations (triangles, quadrilaterals, circles, etc.).
+Each caption was independently evaluated by three annotators with a background in mathematics education at the college level.
+Annotators were asked to check whether every geometric element described in the caption (e.g., points, lines, angles, shapes) was consistent with the diagram rendered from code.
+The task was framed as a binary decision: correct (all elements accurately represented) vs.
+incorrect (at least one element misrepresented or missing).
+All three annotators agreed on every instance, yielding 100% accuracy with full inter-annotator agreement.
+Although the agreement was perfect on this sample, the evaluation was limited to 123 captions.
+Larger-scale or more diverse verification may reveal edge cases where captions misrepresent subtle geometric relations.
+Appendix G Prompts
+After multiple rounds of experimentation, we finalized six prompts for our data construction, as illustrated in Figures 10 to 15.
+Specifically, the prompt in Figure 10 input a textual geometry problem along with its answer and generates a corresponding Python script that renders the geometric image.
+Based on this script, we further generate a positive caption using the prompt shown in Figure 11.
+Directly perturbing the Python script alone to get image often leads to uncontrolled.
+Therefore, as shown in Figure 12, we first generate negative captions by introducing subtle but semantically significant modifications to the positive caption.
+We then use these negative captions to perturb the original Python script, ultimately producing visually plausible but negative images, as shown in Figure 13.
+Since the images are programmatically generated, errors in the scripts are sometimes inevitable.
+To ensure quality, we employ a LLM to automatically verify and correct any syntactic or semantic issues in the generated code by Figure 14.
+For rule-based text negatives, we design specialized prompts to enforce targeted modifications.
+The corresponding prompt is illustrated in Figure 15, which enables controlled perturbations based on observed reasoning errors.
+Figure 10: Prompt for generating a geometric Python script from a given geometry problem and its solution. The resulting script is used to render the corresponding image.
+Figure 11: Prompt for generating a positive geometric caption from a Python script.
+Figure 12: Prompt for generating negative captions by introducing fine-grained semantic modifications to the positive caption.
+Figure 13: Prompt for generating negative geometric images by modifying the original Python script based on the perturbed negative captions.
+Figure 14: Prompt for automatic correction of syntactic or semantic errors in Python scripts using an LLM. This ensures that all generated diagrams are valid and executable.
+Figure 15: Prompt for constructing rule-based text negatives through controlled perturbations of the original caption.
+Appendix H Case Study of Problem Solving
+Figure 16: Comparison of GPT-4o and MMGeoLM in geometric problem-solving.
+MMGeoLM correctly answers the problem using a different solution compared to True Answer, while GPT-4o generates a non-existent element $\triangle ABD$.
+Figure 17: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
+In Figures 16 and 17, we compare the solutions generated by GPT-4o and our MMGeoLM.
+In Figure 16, the red text highlights GPT-4o’s misrecognition of two triangles in the image, leading to an incorrect final result.
+In contrast, MMGeoLM produces a correct solution, albeit through a reasoning process different from the ground-truth answer.
+This highlights MMGeoLM’s capability to generate diverse problem-solving approaches.
+Figure 17 illustrates errors in image-based reasoning for both MMGeoLM and GPT-4o.
+While both models make mistakes, MMGeoLM’s reasoning aligns more closely with the correct solution, demonstrating improved mathematical reasoning.
+Overall, our trained MMGeoLM model demonstrates greater reasoning diversity and enhanced geometric problem-solving performance compared to current mainstream models.